# Try-Mopsa:

### v1

@@ Abstract
Keywords:
Static Analysis Abstract Interpretation Usability Teaching.
@@ 1 Introduction
One remedy to this issue is to provide zero-install ways to try a software, for example by enabling tool usage through a web browser. This article presents Try-Mopsa, a scaled-down version of the Mopsa static analysis platform that runs entirely as a client-side web application. It relies on a responsive interface to support a wide variety of devices (from smartphones to computers), and supports all core features of Mopsa, including relational domains and an interactive engine acting as an abstract debugger. Try-Mopsa is available online [30]. Thanks to its purely client-side implementation, Try-Mopsa scales to a large number of concurrent users.
We provide a brief overview of Mopsa in Section˜2. Then, we describe in Section˜3 how we adapted Mopsa to make it runnable in a web page. Section˜4 provides an overview of the resulting web interface, ˜5 evaluates several features of the implementation, and ˜6 discusses related work.
@@ 2 A Brief Overview of Mopsa
Mopsa is a Modular Open Platform for Static Analysis, rooted within the abstract interpretation framework [6]. It aims at providing a convenient platform for static analysis learners, developers and users. Although Mopsa explores some new perspectives for the design of static analyzers, it is stable and precise enough to be on-par with state-of-the-art academic program analyzers participating to the Software-Verification Competition [2, 33]. [18] describe the core of Mopsa’s principles, and [29, Chapter 3 ] provides an in-depth introduction to Mopsa’s architecture. We briefly describe three features of interest for this article:
@@
Different user interfaces are offered by Mopsa. The automatic interface is the classic one: the analysis runs to completion and then displays the results. The interactive engine lets user navigate the abstract execution of the program, where analysis computations are performed on-the-fly accordingly. This interface acts as a gdb-like abstract debugger, and supports breakpoints, program navigation, printing of abstract states, … The DAP interface provides a debug adapter protocol interface, similar to the interactive engine but allowing interactions with IDEs supporting this protocol. [32, Section 5 ] describes the former two interfaces in more details.
@@ 3.1 Compiling Mopsa to JavaScript
Mopsa depends on the Apron library [17] to support relational numeric domains such as octagons [24] and polyhedra [7]. However, Apron abstract domains are written in C/C++: compiling them to JavaScript, just like our C frontend, would be quite cumbersome. For Try-Mopsa, we chose to replace Apron with the alternative Verified Polyhedra Library (VPL) [3]. Given that some domains of the VPL are pure OCaml implementations, JSOO is able to compile it to JavaScript, enabling relational abstract domains to run within web browsers. We wrote a new VPL wrapper, making it compatible with Mopsa’s interface, and adding support for the fold/expand operators [12] required by Mopsa but not built in the VPL.
@@ 3.2 Handling Web Interactions
Try-Mopsa is split between a toplevel, handling the page interaction with the user and a web worker, handling the actual computations of the analysis. Thanks to this decoupling, the web worker and the page rendering lies within different threads of the web browser, so longer computations do not freeze the browser’s rendering process.
@@
When a user starts an analysis, the toplevel creates a web worker to perform it. The user may choose to interrupt the analysis, in which case the top-level then signals the web worker to stop.
@@
The interactive engine of Mopsa follows an interaction loop by awaiting user input (in a synchronous, blocking fashion) and providing corresponding results. However, web interfaces typically rely on asynchronous processes. We explored options such as implementing an asynchronous interactive engine. This would have required the decoupling of the interactive interface from the analysis computations, which would have been a large implementation effort introducing breaking changes. In the end, we chose to rely on the sync-message [27] JavaScript library enabling synchronous communication between the web worker to the toplevel, should some user input be required.
@@ 4 Try-Mopsa: Interface Overview
This loop may stop non-deterministically at any iteration (line 6).
@@
s among $S=\{\text{"}a\text{"},\text{"}ab\text{"},\text{"}abc\text{"},\ldots,\text{"}abc\cdots{}xyz\text{"}\}$. The assertion at line 14 checks a simpler property: that s belongs $S$ but up to any permutation of the characters.
@@
˜4 shows the configuration editor, letting users specify their choice of abstract domains and combinators. Similarly to the code editor, users can also load a configuration from a pre-defined list. We comment on the loaded configuration, string_product_relational.json. This configuration features three notable abstract domains, working together. The string length domain handles ghost variables representing the length of each string. The string summarization handles ghost variables representing a summary of the contents of each string, through the ASCII codes of the contained characters. These two domains are defined as a reduced product as they infer complementary pieces of information. Finally, a relational numerical domain is enabled. It handles the constraints passed by the aforementioned string domains, which allows it to infer relations between those different ghost variables, and program variables. The result of analyzing str_alphabet2.u using string_product_relation-al.json is displayed at the bottom of Figure˜3. In this case, Try-Mopsa is able to prove the complex assertion at line 18, which mixes information about the contents and the length of string s. Let us briefly comment on the abstract state inferred after the loop and printed at lines 1-9 of the analysis result in Figure˜3. We can see that the relational domain inferred equality between the length of the string s and the integer variable i. We can further notice an interesting relation between the contents of s ($\text{ord}(s)$) and its length: $\text{'}a\text{'}\leq\text{ord}(s)\leq\text{'}a\text{'}+\text{len}(s)-1$. This invariant has been obtained thanks to the cooperation between the string abstract domains and the relational abstract domain. Note that the second inequality is exactly the invariant proved by Mopsa at line 18.
@@
˜5 displays various analysis options to tweak the analysis. As we mentioned in the previous section, these options are dynamically generated by the toplevel, depending on which abstract domains are activated. In this example, the first five options are defined at the framework level, while the last four options customize the behavior of the loop iterator. Note that the options argument can be selected through different interfaces depending on the expected output (e.g., choice among fixed list, integer, string). To further improve its usability, Try-Mopsa can also be used as a progressive web app, meaning that it can be installed locally on a device and run without relying on the network. In addition, a share button (in the top-right corner of Figure˜3) lets users share their current program, configuration and option choices by encoding it into a URL. This can be useful for example to simplify practical sessions where no manual loading is required.
@@ 5 Implementation Discussion
We argue that the implementation of Try-Mopsa is maintainable with respect to the standard implementation of Mopsa. Indeed, Try-Mopsa does not introduce breaking changes, meaning future development can be shared in the same repository. In addition, most components of Mopsa were reused as-is, thanks in particular to the reuse of the same input-output loop as a terminal user of Mopsa would. The only major component that has been added is the support for relational abstract domains through the VPL library. This implementation provides adequate results on the toy examples we provide along with Try-Mopsa.
Table 1: Comparison of analysis times with relational abstract domains enabled. The mean and standard deviation have been computed over 10 runs.
@@
The goal of Try-Mopsa is to ease tool demonstrations and testing without installation, where raw analysis performance is not a priority. Nevertheless, we compared the analysis times between the natively generated executable, and the same code running in Firefox. We ran this comparison on the toy, default examples of Try-Mopsa, with configurations using relational abstract domains. We show in ˜1 the nine most significant programs to analyze. On average, Try-Mopsa is five times slower than native binary execution. While this is a noticeable slowdown, we believe the analysis times – below one second – make Try-Mopsa still usable for demonstration or teaching purposes. Future work could focus on using JSOO’s recent WebAssembly code generation features to obtain more efficient code.
@@ 6 Related Work
In previous works, developers relied on server-side computations to provide web interfaces for static analyzers relying on relational numerical domains. This includes for example Interproc [15, 16], Banal [22, 25] and FuncTion [44, 45]. While client-server implementations require less work to adapt a static analyzer to the web, these implementations raise more security concerns and can be less scalable in the number of users. In addition, these implementations need to maintain a compatible server service over the years. This proves to be difficult: at the time of writing, the majority of these approaches are no longer operational. More recently, pure client-side web interfaces have been developed for static analyzers. These scale much better with the number of concurrent users, and their “only” dependency is a web browser – one of the most widespread pieces of software. [19, 20] provide a demo web interface for their Salto analyzer of OCaml code, following previous demonstrators accompanying works around the static analysis of functional languages [35, 36]. In their BINSEC [9] tutorial at PLDI 2025, [41] provide a web tutorial where users can choose between running BINSEC in CLI or running analysis scenarios directly in their browser. To the best of our knowledge, Try-Mopsa is the only web interface supporting the analysis of multiple languages, relational domains, and an interactive exploration of the analysis. [37] describe how they use their static analysis platform LiSA within their static analysis courses. Students are tasked with installing LiSA, and implementing several simple abstract domains. Try-Mopsa aims at providing a zero-install static analyzer relying on complex abstract domains – such as polyhedra – to illustrate how such a tool works. Evaluating the impact of Try-Mopsa on real teaching cohorts is left as future work. Following studies about developer use of static analyzers [5, 11], there are quite a few works discussing interfaces for static analyzers. MagpieBridge [21] aims at simplifying the display of analysis results within IDEs supporting the Language-Server Protocol. Several interfaces have also been developed to debug static analyzers [10], provide an abstract debugger [13], or mix concrete and abstract debuggers [28]. Outside of the static analysis community, [4] describe the technical components requiring to run an OCaml MOOC who attracted thousands of learners. It relies in particular on a pure client-side development environment avoiding installation woes students could otherwise face. [1] also provide a pure JavaScript version for the Rocq proof assistant, which aims at improving literate programming formatting and reducing installation burdens, in an educational context.
@@ 7 Conclusion
This article introduced Try-Mopsa, a pure, client-side implementation of Mopsa running in the browser or as a progressive web app. Try-Mopsa supports all the main core components of Mopsa, including support for relational numerical domains and its abstract debugger. Try-Mopsa is designed to work on a wide range of devices, from smartphones to desktop, thanks to a responsive design and continuous testing of the support of popular browsers and various resolutions. We envision Try-Mopsa to be a convenient tool to demonstrate static analysis capabilities, for teaching purposes or more generally to attract new users.
@@ Acknowledgments.
We are grateful to Aymane Ismail for prototyping an early version of Try-Mopsa; and to Alexandre Maréchal for his help in using the VPL. We thank Aymeric Fromherz, Louis Gesbert, Vincent Botbol, Louis Rustenholz, Antoine Miné, Abdelraouf Ouadjaout and the whole Mopsa team for their helpful discussion and comments on this work.

### v2

# UMR 9189 CRIStAL, F-59000 Lille, France

@@ 1 Introduction
One remedy to this issue is to provide zero-install ways to try a software, for example by enabling tool usage through a web browser. This article presents Try-Mopsa, a scaled-down version of the Mopsa static analysis platform that runs entirely as a client-side web application. It relies on a responsive interface to support a wide variety of devices (from smartphones to computers), and supports all core features of Mopsa, including relational domains and an interactive engine acting as an abstract debugger. Try-Mopsa is available online [38]. Thanks to its purely client-side implementation, Try-Mopsa is unaffected by the scalability issues of server-side implementations serving a large number of concurrent users.
We provide a brief overview of Mopsa in Section˜2. Then, we describe in Section˜3 how we adapted Mopsa to make it runnable in a web page. Section˜4 provides an overview of the resulting web interface, Section˜5 evaluates several features of the implementation, and Section˜6 discusses related work.
@@ 2 A Brief Overview of Mopsa
Mopsa is a Modular Open Platform for Static Analysis, rooted within the abstract interpretation framework [10]. It aims at providing a convenient platform for static analysis learners, developers and users. Although Mopsa explores some new perspectives for the design of static analyzers, it is stable and precise enough to be on-par with state-of-the-art academic program analyzers participating to the Software-Verification Competition [4, 42]. [24] describe the core of Mopsa’s principles, and [37, Chapter 3 ] provides an in-depth introduction to Mopsa’s architecture. We briefly describe three features of interest for this article:
@@
Different user interfaces are offered by Mopsa. The automatic interface is the classic one: the analysis runs to completion and then displays the results. The interactive engine lets user navigate the abstract execution of the program, where analysis computations are performed on-the-fly accordingly. This interface acts as a gdb-like abstract debugger, and supports breakpoints, program navigation, printing of abstract states, … The DAP interface provides a debug adapter protocol interface, similar to the interactive engine but allowing interactions with IDEs supporting this protocol. [41, Section 5 ] describe the former two interfaces in more details.
@@ 3.1 Compiling Mopsa to JavaScript
Mopsa depends on the Apron library [23] to support relational numeric domains such as octagons [31] and polyhedra [11]. However, Apron abstract domains are written in C/C++: compiling them to JavaScript, just like our C frontend, would be quite cumbersome. For Try-Mopsa, we chose to replace Apron with the alternative Verified Polyhedra Library (VPL) [5], which provides different implementations of polyhedra. Given that some domains of the VPL are pure OCaml implementations, JSOO is able to compile it to JavaScript, enabling relational abstract domains to run within web browsers. We wrote a new VPL wrapper, making it compatible with Mopsa’s interface, and adding support for the fold/expand operators [18] required by Mopsa but not built in the VPL.
@@ 3.2 Handling Web Interactions
Try-Mopsa is split between a toplevel, handling the page interaction with the user and a web worker, handling the actual computations of the analysis. Thanks to this decoupling, the web worker and the page rendering lie within different threads of the web browser, so longer computations do not freeze the browser’s rendering process.
@@
When a user starts an analysis, the toplevel creates a web worker to perform it. The user may choose to interrupt the analysis, in which case the toplevel then signals the web worker to stop.
@@
The interactive engine of Mopsa follows an interaction loop by awaiting user input (in a synchronous, blocking fashion) and providing corresponding results. However, web interfaces typically rely on asynchronous processes. We explored options such as implementing an asynchronous interactive engine. This would have required the decoupling of the interactive interface from the analysis computations, which would have been a large implementation effort introducing breaking changes. In the end, we chose to rely on the sync-message [35] JavaScript library enabling synchronous communication between the web worker to the toplevel, should some user input be required. Thanks to this embedding, all user-interface features of Mopsa are supported in Try-Mopsa.
@@ 4 Try-Mopsa: Interface Overview
This loop may stop non-deterministically at any iteration (line 10).
@@
s among $S=\{\text{"}a\text{"},\text{"}ab\text{"},\text{"}abc\text{"},\ldots,\text{"}abc\cdots{}xyz\text{"}\}$. The assertion at line 18 checks a simpler property: that s belongs to $S$ but up to any permutation of the characters.
@@
Figure˜4 shows the configuration editor, letting users specify their choice of abstract domains and combinators. Similarly to the code editor, users can also load a configuration from a pre-defined list. We comment on the loaded configuration, string_product_relational.json. This configuration features three notable abstract domains, working together. The string length domain handles ghost variables representing the length of each string. The string summarization handles ghost variables representing a summary of the contents of each string, through the ASCII codes of the contained characters. These two domains are defined as a reduced product as they infer complementary pieces of information [37]. Finally, a relational numerical domain is enabled. It handles the constraints passed by the aforementioned string domains, which allows it to infer relations between those different ghost variables, and program variables.
The result of analyzing str_alphabet2.u using string_product_relation-al.json is displayed at the bottom of Figure˜3. In this case, Try-Mopsa is able to prove the complex assertion at line 18, which mixes information about the contents and the length of string s. Let us briefly comment on the abstract state inferred after the loop and printed at lines 1-9 of the analysis result in Figure˜3. We can see that the relational domain inferred equality between the length of the string s and the integer variable i. We can further notice an interesting relation between the contents of s ($\text{ord}(s)$) and its length: $\text{ord}(s)\leq\text{'}a\text{'}+\text{len}(s)-1$. This invariant has been obtained thanks to the cooperation between the string abstract domains and the relational abstract domain. Note that the second inequality is exactly the invariant proved by Mopsa at line 18.
@@
Figure˜5 displays various options to tweak the analysis. As we mentioned in the previous section, these options are dynamically generated by the toplevel, depending on which abstract domains are activated. In this example, the first five options are defined at the framework level, while the last four options customize the behavior of the loop iterator. Note that the option’s argument can be selected through different interfaces depending on the expected output (e.g., choice among fixed list, integer, string).
To further improve its usability, Try-Mopsa can also be used as a progressive web app, meaning that it can be installed locally on a device and run without relying on the network. In addition, a share button (in the top-right corner of Figure˜3) lets users share their current program, configuration and option choices by encoding it into a URL. This can be useful for example to simplify practical sessions where no manual loading is required.
@@ 5 Implementation Discussion
We argue that the implementation of Try-Mopsa is maintainable with respect to the standard implementation of Mopsa. Indeed, Try-Mopsa does not introduce breaking changes, meaning future development can be shared in the same repository.
In addition, most components of Mopsa were reused as-is, thanks in particular to the reuse of the same input-output loop as a terminal user of Mopsa would. The only major component that has been added is the support for relational abstract domains through the VPL library. This implementation provides adequate results on the toy examples we provide along with Try-Mopsa.
Table 1: Preliminary evaluation of analysis times with relational abstract domains enabled. The mean and standard deviation have been computed over 10 runs.
@@
The goal of Try-Mopsa is to ease tool demonstrations and testing without installation, where raw analysis performance is not a priority. Nevertheless, we ran a preliminary evaluation comparing the analysis times between the natively generated executable, and the same code running in Firefox. We ran this comparison on the toy, default examples of Try-Mopsa, with configurations using relational abstract domains. We show in Table˜1 the nine most significant programs to analyze. On average, Try-Mopsa is five times slower than native binary execution. While this is a noticeable slowdown, we believe the analysis times – below one second – make Try-Mopsa still usable for demonstration or teaching purposes. Future work could focus on using JSOO’s recent WebAssembly code generation features to obtain more efficient code.
@@ 6 Related Work
In previous works, developers relied on server-side computations to provide web interfaces for static analyzers relying on relational numerical domains. This includes for example Interproc [21, 22], Banal [29, 32] and FuncTion [54, 55]. While client-server implementations require less work to adapt a static analyzer to the web, these implementations raise more security concerns and can be less scalable in the number of users. In addition, these implementations need to maintain a compatible server service over the years. This proves to be difficult: at the time of writing, the majority of these approaches are no longer operational.
More recently, pure client-side web interfaces have been developed for static analyzers. These scale much better with the number of concurrent users, and their “only” dependency is a web browser – one of the most widespread pieces of software. [26, 27] provide a demo web interface for their Salto analyzer of OCaml code, following previous demonstrators accompanying works around the static analysis of functional languages [44, 45]. [16, 46] provide a WebAssembly version of Ciao Prolog, embedding the CiaoPP analyzer, which also supports interactive user inputs. In their BINSEC tutorial at PLDI 2025, [51] provide a web tutorial where users can choose between running BINSEC [13] in CLI or running analysis scenarios directly in their browser. To the best of our knowledge, Try-Mopsa is the only web interface supporting the analysis of multiple languages (our toy imperative language and Python), numerical relational domains, and an interactive exploration of the analysis.
Note that another lightweight approach for trying out some new software is the use of Docker containers. We actually provide Docker containers with each release of Mopsa [34]; these containers supports all features of Mopsa (e.g., including the C analysis).
[47] describe how they use their static analysis platform LiSA within their static analysis courses. Students are tasked with installing LiSA, and implementing several simple abstract domains. Try-Mopsa aims at providing a zero-install static analyzer relying on complex abstract domains – such as polyhedra – to illustrate how such a tool works. Evaluating the impact of Try-Mopsa on real teaching cohorts is left as future work.
Following studies about developer use of static analyzers [7, 15], there are quite a few works discussing interfaces for static analyzers. MagpieBridge [28] aims at simplifying the display of analysis results within IDEs supporting the Language-Server Protocol. Several interfaces have also been developed to debug static analyzers [14], provide an abstract debugger [19], or mix concrete and abstract debuggers [36]. [9] recently introduced a new graphical interface for Frama-C users.
Outside of the abstract interpretation community, [25] provide a unit testing as a service platform. Both the Alt-Ergo SMT Solver [8] and the Why3 deductive verification platform [17] leverage JSOO to provide zero-install web versions [1, 3]. [6] describe the technical components required to run an OCaml MOOC (Massive Open Online Course) who attracted thousands of learners. It relies in particular on a pure client-side development environment avoiding installation woes students could otherwise face. [2] also provide a pure JavaScript version for the Rocq proof assistant, which aims at improving literate programming formatting and reducing installation burdens, in an educational context.
@@ 7 Conclusion
This article introduced Try-Mopsa, a pure client-side implementation of Mopsa running in the browser or as a progressive web app. Try-Mopsa supports all the main core components of Mopsa, including support for relational numerical domains and its abstract debugger. Try-Mopsa is designed to work on a wide range of devices, from smartphones to desktops, thanks to a responsive design and continuous testing of the support of popular browsers and various resolutions. We envision Try-Mopsa to be a convenient tool to demonstrate static analysis capabilities, for teaching purposes or more generally to attract new users.
Data Availability Statement.
This paper is accompanied by an artefact [39], which has been peer-reviewed as “available, functional and reusable”.
@@ Acknowledgments.
We thank the anonymous VMCAI reviewers and the participants of Dagstuhl Seminar #25421 for their valuable feedback. We are grateful to Aymane Ismail for prototyping an early version of Try-Mopsa; and to Alexandre Maréchal for his help in using the VPL. We thank Aymeric Fromherz, Louis Gesbert, Vincent Botbol, Louis Rustenholz, Antoine Miné, Abdelraouf Ouadjaout and the whole Mopsa team for their helpful discussion and comments on this work.